If 490 nm light falls on a slit 0.0430 mm wide, what is the full angular width of the central diffraction peak?
Answer #1
Concepts and reason

The required concept to solve the problem is the single slit diffraction experiment.

First, from the condition for diffraction maxima, find the angle between the normal line between the screen and slit and the wave ray to a point on the screen.

Then, using the angle between the normal line between the screen, slit, and the wave ray to a point on the screen and the angular width, find the angular width of the central diffraction peak.

Fundamentals

A light can act both as a particle and as a wave. It can bend around the objects like a wave. This is the concept behind the single slit diffraction. Due to interference, the waves are added together constructively, and then, a maxima is formed. If the waves are added together destructively, then a minima is formed.

The equation for diffraction maxima is,

dsinθ=mλd\sin \theta = m\lambda

Here, dd is the width of the slit, θ\theta is the angle between the normal line between the screenand slit, and the wave ray to a point on the screen, mm is the order, and λ\lambda is the wavelength.

The angular width is,

Φ=2θ\Phi = 2\theta

The condition for diffraction maxima is,

dsinθ=mλd\sin \theta = m\lambda

Rearrange the above equation in terms of θ\theta .

sinθ=mλdθ=sin1(mλd)\begin{array}{c}\\\sin \theta = \frac{{m\lambda }}{d}\\\\\theta = {\sin ^{ - 1}}\left( {\frac{{m\lambda }}{d}} \right)\\\end{array}

Substitute 490nm490\,{\rm{nm}} for λ\lambda , 11 for mm and 0.0430mm0.0430\,{\rm{mm}} for dd to find θ\theta .

θ=sin1[(1)((490nm)(109m1nm))(0.0430mm)(103m1mm)]=sin1[(490×109m)(0.0430×103m)]=sin1(0.01139)=0.653\begin{array}{c}\\\theta = {\sin ^{ - 1}}\left[ {\frac{{\left( 1 \right)\left( {\left( {490\,{\rm{nm}}} \right)\left( {\frac{{{{10}^{ - 9}}\,{\rm{m}}}}{{1\,{\rm{nm}}}}} \right)} \right)}}{{\left( {0.0430\,{\rm{mm}}} \right)\left( {\frac{{{{10}^{ - 3}}\,{\rm{m}}}}{{1\,{\rm{mm}}}}} \right)}}} \right]\\\\ = {\sin ^{ - 1}}\left[ {\frac{{\left( {490 \times {{10}^{ - 9}}\,{\rm{m}}} \right)}}{{\left( {0.0430 \times {{10}^{ - 3}}\,{\rm{m}}} \right)}}} \right]\\\\ = {\sin ^{ - 1}}\left( {0.01139} \right)\\\\ = 0.653^\circ \\\end{array}

The equation for the angular widthis,

Φ=2θ\Phi = 2\theta

Substitute 0.6530.653^\circ for θ\theta .

Φ=2(0.653)=1.306\begin{array}{c}\\\Phi = 2\left( {0.653^\circ } \right)\\\\ = 1.306^\circ \\\end{array}

Ans:

Thus, the full angular width of the central diffraction peak is 1.3061.306^\circ .